OpenStudy (anonymous):

for the function u=t^3/2 show that the derivative does not exist at t=0 but that the right-hand derivative exist at that point

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Are you expected to use the limit definition for this? Or can you use the friendly rules of differentiation?

OpenStudy (anonymous):

im expected to use the limit of definition

zepdrix (zepdrix):

\[\large u(t)=t^{3/2}\] Using the Limit Definition of the Derivative:\[\large u'(t)=\lim_{h \to 0}\frac{(t+h)^{3/2}-t^{3/2}}{h}\] And from here, let's seeee.

zepdrix (zepdrix):

Multiplying the top and bottom by the `conjugate of the top` gives us,\[\large u'(t)=\lim_{h \to 0}\frac{(t+h)^{3/2}-t^{3/2}}{h}\left(\frac{(t+h)^{3/2}+t^{3/2}}{(t+h)^{3/2}+t^{3/2}}\right)\] \[\large u'(t)=\lim_{h \to 0}\frac{(t+h)^3-t^3}{h\left[(t+h)^{3/2}+t^{3/2}\right]}\]

zepdrix (zepdrix):

Understand that step? It's a little tricky :o

zepdrix (zepdrix):

Err hmm, maybe we don't need to go into all of that.. maybe we can simply do this,

zepdrix (zepdrix):

no, i think we're on the right track..

OpenStudy (anonymous):

the multiplication of the conjugate top is confusing can you elaborate further please :)

zepdrix (zepdrix):

Normally when we multiply conjugates it's a little more straight forward, like this:\[\large (a-b)(a+b) \qquad=\qquad a^2-b^2\] But in our case we have something like this:\[\large \left(a^{3/2}-b^{3/2}\right)\left(a^{3/2}+b^{3/2}\right) \qquad=\qquad \left(a^{3/2}\right)^2-\left(b^{3/2}\right)^2\] When we have an exponent with another exponent being applied like this, our rules of exponents tell us to `multiply the exponents`\[\large \frac{3}{2}\cdot 2 \qquad=\qquad 3\]So we're ending up with,\[\large a^3-b^3\]

zepdrix (zepdrix):

too much? did your head esplode?